Horn Company is considering the purchase of a new machine for $108,000.
The machine would replace an old piece of equipment that costs $41,830
per year to operate. The new machine would cost $25,720 per year to
operate. The old machine currently in use can be sold for $9,500 if
the new machine is purchased. The new machine would have a useful life
of ten years with a $6,000 salvage value.

Calculate the accounting rate of return on the machine that Horn Company
is considering buying. Enter your answer as a number without the % symbol.
For example, if your answer is 10%, simply enter 10 as your answer.
Answer #1
Annual Depreciation on new machine 10200 =(108000-6000)/10
Annual savings in Operating costs 16110 =41830-25720
Less: Annual Depreciation on new machine 10200
Annual net income 5910
Annual net income 5910
Divide by Net initial investment 98500 =108000-9500
Accounting rate of return 6